OpenStudy (anonymous):

PLEASEE PLEASEEE ITS A TEST GRADE ;( the volumes of two similar rectangualr solids are 125 cm^3 and 8 cm^3. Find the ratio of their base perimeter. SHOW WORK. pleasee

OpenStudy (e.mccormick):

How did you start and how far did you get, or what is the part you do not get?

OpenStudy (anonymous):

I'm not sure how to it at all

OpenStudy (e.mccormick):

OK. Lets start with a few terms. Similar means they have the same ratios to the sides. So they are not the same, but they have the same basic shape.

OpenStudy (anonymous):

yes, i understand that

OpenStudy (e.mccormick):

OK. And you know the volume is height*width*depth, right? So you have these two bits of information.

OpenStudy (e.mccormick):

The ratios work in a similar way. Lets say B represents measure of the big one and S the small. If I do a ratio like this for the volumes:\[\frac{B^3}{S^3}\]The key point becomes the fact that they are cubic. That is volume. Area is square. Now the ratios are related, so the ratio of the square area of something with \(\frac{B^3}{S^3}\) is relationship in the form of:\[\frac{B^2}{S^2}\]

OpenStudy (e.mccormick):

Let me show this with two cubes. That will show the general idea. These will not be your rectangles, but they should get the point across.